The Wrightsville Rakes Company produces various types of yard rakes. It has calculated the average cost per unit of a production level of 7,900 rakes to be $12. If $22,900 of the costs are fixed, and the plant manager uses the cost equation to predict total cost, her forcast for 8,900 rakes will be (Round any intermediary calculation to the nearest cent.)

A) $94,800

B) $12,000

C) 103,890

D) 106, 800
Answer #1

The answer is C) 103,890


7,900 * $12 = 94,800

94,800 - 22,900 = 71,900

71,900 / 7,900 = 9.1

(9.1*8900)+22,900=103,890